OpenStudy (anonymous):

My answer is C What is the 41st term of the arithmetic sequence where a1 = 17 and a15 = -39 ? (3 points) Answer -147 -143 -139 -135

OpenStudy (anonymous):

-39 = 16 + (15 - 1) d d = -3.92 a41 = 16 + (41 -1 ) x -3.92 = 16 + 40 x ( -3.9) = -140

OpenStudy (anonymous):

yep i guess your answer is right

OpenStudy (anonymous):

Hmmm I wonder how you got a number higher

hartnn (hartnn):

17-160 = -143

OpenStudy (anonymous):

it depends on the value of d, common difference I used 3.9 in my calculations

hartnn (hartnn):

d=-4

hartnn (hartnn):

-39 = 17 + (15 - 1) d d=-4

OpenStudy (anonymous):

Ohhhhhh hartnn is right im wrong a1 = 17 i used 16..opss

OpenStudy (anonymous):

Thank you